OpenStudy (anonymous):

got a long one f(x) = x+5 and g(x) = x+8 find the following (f+g)(x) then (f-g)(x) and then (fg)(x) last one f/g)(x)

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x) similar for (f-g)(x) (fg)(x)=f(x)g(x) (f/g)(x)=f(x)/g(x)

OpenStudy (anonymous):

? thats not the answers?

OpenStudy (anonymous):

u r to substitute the values only

OpenStudy (anonymous):

do u know how to do it?

OpenStudy (anonymous):

no? Im stuck I keep getting different answers

OpenStudy (anonymous):

ok (f+g)(x)=f(x)+g(x)=x+5+x+8=2x+13

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

can u do the rest?

OpenStudy (anonymous):

I can do the next step,,, how would I do the (fg)(x) one?

OpenStudy (anonymous):

(fg)(x)=f(x)g(x)=(x+5)(x+8)=x^2+13x+40

OpenStudy (anonymous):

the (f-g) would be 2x-3 for the answer right?

OpenStudy (anonymous):

it would be -3 only, x is cancelled out

OpenStudy (anonymous):

oh..crap k

OpenStudy (anonymous):

did u get the product ?

OpenStudy (anonymous):

last one would be x+5/x+8

OpenStudy (anonymous):

right

OpenStudy (anonymous):

sweet ... think I get it now thank you :)

OpenStudy (anonymous):

:)